lightfoot inc, a software development firm has a stock outstanding as follows 15000 shares of cumulative preferred 3% stock, $20 par, and 19000 shares of $125 par common. During its first four years of operations, the following amounts were distributed as dividends: first year, 3,450; second year, $5,700; third year, $34,570; fourth year, $65810.
calculate the dividends per share on each class of stock for each of the four years. round all answers to two decimal places. if no dividends are paid in a given year, enter 0
Answer #1
year preferred share dividend per share common stock dividend per share
1 $0.23 nil
2 $0.38 nil
3 $1.19 0.88
4 $0.60 2.99

working:

annual preferred dividend payment = (15,000*$20*3%)=>$9,000.

first year dividend = $3,450

preferred dividend per share = 3450/15000=>$0.23

first year dividend carried forward next year = 9,000-3450

=>5,550

second year dividend to be paid = $9,000 + 5,550 carried forward =>14,550.

second year dividend paid = $5,700.

preferred dividend per share =>$5700/15,000=>$0.38

second year dividend carried forward = 14,550-5,700 =>8,850.

third year dividend to be paid = $9,000 + 8,850 =>$17,850.

third year total dividend paid = 34,570

=>preferred dividend paid = 17,850

preferred dividend per share = 17,850/15,000=>$1.19

common stock dividend paid = 34,570-17,850

=>$16,720.

common stock dividend per share =16,720/19,000=>$0.88

fourth year dividend paid = $65,810 in total

preferred dividend paid = $9,000

per share preferred dividend = $9000/15,000 =>$0.60

ccmmon stock dividend =65,810-9,000

=>$56,810

per share common stock = $56,810/19,000

=>$2.99.